# DistRL: An Asynchronous Distributed Reinforcement Learning Framework for On-Device Control Agents

### v3

@@ 5.2 Distributed Prioritized Experience Replay (DPER)
To improve sample efficiency, we employ Distributed Prioritized Experience Replay (DPER). For each trajectory $\tau=\{(s_{t},a_{t},r_{t},s_{t+1})\}_{t=0}^{H}$, we compute the priority $p(\tau)$ as: $p(\tau)=w_{1}\overline{|\delta|}+w_{2}\overline{\rho}+w_{3}\overline{\mathbb{H}},$ where $\overline{|\delta|}$ is the average absolute temporal-difference (TD) error over the trajectory, calculated as $\delta_{t}=r_{t}+\gamma V(s_{t+1})-V(s_{t})$; $\overline{\rho}$ is the average importance sampling ratio $\rho_{t}$; and $\overline{\mathbb{H}}$ is the average policy entropy, $\mathbb{H}_{t}=-\log\pi(a_{t}|s_{t})$, encouraging exploration by encouraging policy uncertainty, thus avoiding early convergence to suboptimal policies during training in dynamic environments. The weights $w_{1}$, $w_{2}$, and $w_{3}$ balance the contributions of each component, which is selected by grid-search. Trajectories with higher priorities are replayed more frequently, focusing learning on the most informative experiences. Priorities are periodically updated based on the latest policy, recalculating them to focus learning on the most informative experiences, ensuring continual adaptation to evolving behavior policies. Details can be found in Appendix A.4.
5.3 DistRL Pipeline Implementation
As illustrated in Figure 4, DistRL adopts a distributed asynchronous setup where multiple worker agents generate trajectories under the behavior policy $\mu$ and send them to a central learner. The trajectory reward is computed using the Monte Carlo estimate:
$L(V_{\text{traj}})=-\mathbb{E}_{\nu}\left[r(s_{H},a_{H})\log V_{\text{traj}}(s_{H},a_{H})+\left(1-r(s_{H},a_{H})\right)\log\left(1-V_{\text{traj}}(s_{H},a_{H})\right)\right].$ (3)
The actor is updated using policy gradients based on advantage estimates, and enhanced Retrace corrections are applied for off-policy learning. This process is distributed asynchronously across worker nodes, ensuring efficient fine-tuning in environments with sparse rewards and distributed delays.
@@ 6 Experiments
To evaluate the performance of DistRL on challenging Android device control tasks, we conducted extensive experiments. Our primary goal is to determine whether DistRL can produce agents that effectively learn from autonomous online interaction. We describe the experimental environment in Section 6.1, the baseline and benchmarks in Section 6.2, training performance in Section 6.3, and on-device task evaluations in Section 6.4. Additionally, we present ablation studies on our approach’s components in Section 6.5.
@@ 6.1 Evaluation Environment
Our evaluation environment consists of a host learner with 4 NVIDIA V100 GPUs for intensive policy training and two worker machines with 8 NVIDIA Tesla T4 GPUs and 96 vCPUs each, supporting parallel emulation. This setup leverages 192 vCPUs to run multiple emulators concurrently, enabling scalable distributed reinforcement learning experiments. The worker machines handle inference and data collection, asynchronously communicating with the host learner to exchange trajectories and updated model weights. This configuration allows us to assess DistRL’s scalability and performance in a realistic large-scale distributed environment.
@@ 6.3 Training Performance
6.4 On-Device End-to-end Agent Performance Evaluation
@@ A.1 Case Study
DistRL successfully completes the task by efficiently navigating through the app’s interface. It opens the correct messaging app and enters the "New Conversation" screen without getting stuck.
@@
GPT-4V, though not specifically trained for app-specific tasks, performs well in this scenario due to its generalization capability. It opens the correct messaging app and navigates to the "New Conversation" screen successfully. GPT-4V is more flexible and suitable for simpler, general-purpose tasks. However, this general-purpose approach may not scale well for more complex tasks where app-specific expertise and interaction nuances are required.
@@
Figure 9 shows a case study on the web shopping task.
DistRL demonstrates relatively smooth and fluid operations, progressing through the steps without hesitation. While it ultimately encounters early termination due to reaching the step limit (horizon), it performs each step with clear transitions and effectively navigates through the sequence. DistRL shows strong task comprehension and adaptation throughout, but its misunderstanding on the task requirement prevents it from fully completing the task. This behavior emphasizes the efficiency of DistRL’s operations and its capacity to generalize across unseen web shopping tasks, even though the task is terminated early.
@@ A.4.1 Limitations of Traditional Methods
Unlike model-free approaches used in applications such as Sun et al. (a; b), Sun & Wang (2022), Wang et al. (2018), Wang & Shi (2021), Wang & Yoneki (2024), Wang et al. (2024), Li et al. (2024b), Wang & Yoneki , Hadian et al. (2020), our RL algorithm for on-mobile training tasks requires customization and adaptation to meet specific task demands.On-policy algorithms such as Proximal Policy Optimization (PPO (Schulman et al., 2017)) and Advantage Actor-Critic (A2C (Mnih, 2016)) require synchronous data collection and updates, leading to inefficiencies in distributed and large-scale environments due to low sample efficiency and synchronization delays.
@@ A.4.2 Implementation Details
where $\alpha=0.5$ controls the extent to which prioritization is applied. A value of $\alpha=0.5$ provides a balance between uniform sampling (when $\alpha=0$) and full prioritization (when $\alpha=1$), allowing the model to benefit from both the prioritization of informative trajectories and a degree of randomness. This ensures that less prioritized but potentially useful experiences still have a chance to be replayed, helping prevent overfitting to a narrow subset of the replay buffer.
@@ A.5.1 Baseline Methods
We evaluate proprietary vision-language models (VLMs), GPT-4V (OpenAI et al., 2024) and Gemini 1.5 Pro (Reid et al., 2024), using the AppAgent framework. By applying the prompt from  (Yang et al., 2023b), we enable these models to interact effectively with the environment.
We assessed the AppAgent (Yang et al., 2023b) in a augmented prompting setting, where the agent explores the environment and gathers experience ahead of inference phase. This collected experience is appended to the test-time prompt, enhancing the model’s decision-making capabilities. Unlike learning-based approaches, these methods rely on advanced prompting strategies to accomplish tasks without extensive training. Additionally, our framework DistRL integrates GPT-2 (Radford et al., 2019) for policy learning and ROBERTA-base (Liu, 2019) for the critic model We adopt the same baseline agent (AutoUI-driven agent) and VLMs models as used in DigiRL to validate the advantages of our fine-tuning framework and algorithm

### v4

@@ 5.3 Distributed Prioritized Experience Replay (DPER)
To improve sample efficiency, we employ Distributed Prioritized Experience Replay (DPER). For each trajectory $\tau=\{(s_{t},a_{t},r_{t},s_{t+1})\}_{t=0}^{H}$, we compute the priority $p(\tau)$ as: $p(\tau)=w_{1}\overline{|\delta|}+w_{2}\overline{\rho}+w_{3}\overline{\mathbb{H}},$ where $\overline{|\delta|}$ is the average absolute temporal-difference (TD) error over the trajectory, calculated as $\delta_{t}=r_{t}+\gamma V(s_{t+1})-V(s_{t})$; $\overline{\rho}$ is the average importance sampling ratio $\rho_{t}$; and $\overline{\mathbb{H}}$ is the average policy entropy, $\mathbb{H}_{t}=-\log\pi(a_{t}|s_{t})$, encouraging exploration by encouraging policy uncertainty, thus avoiding early convergence to suboptimal policies during training in dynamic environments. The weights $w_{1}$, $w_{2}$, and $w_{3}$ balance the contributions of each component, which is selected by grid-search (See Appendix A.4.3). Trajectories with higher priorities are replayed more frequently, focusing learning on the most informative experiences. Priorities are periodically updated based on the latest policy, recalculating them to focus learning on the most informative experiences, ensuring continual adaptation to evolving behavior policies.
@@ 6 Experiments
To evaluate the performance of DistRL on challenging Android device control tasks, we conducted extensive experiments. Our primary goal is to determine whether DistRL can produce agents that effectively learn from autonomous online interaction. We present the experimental environment in Section 6.1, the baseline and benchmarks in Section 6.2, validation of our evaluator in Section 6.3, training performance in Section 6.4, and on-device task evaluations in Section 6.5. Additionally, we present ablation studies on our approach’s components in Section 6.6.
@@ 6.1 Evaluation Environment
Our evaluation environment consists of a host learner with 4 NVIDIA V100 GPUs for intensive policy training and two worker machines with 8 NVIDIA Tesla T4 GPUs and 96 vCPUs each, supporting parallel emulation. This setup leverages 192 vCPUs to run multiple emulators concurrently, enabling scalable distributed reinforcement learning experiments.
@@ 6.2 Benchmarks and Baseline Methods
6.3 Validation of our Evaluator
In our experiments, we used the last screenshot along with the last two actions as input to the VLM Evaluator (Gemini) for prompting. This approach provides a concise yet informative context for the evaluator. As shown in Figure 5, when testing different policy models on General subsets from AiTW, incorporating additional context, such as longer trajectory information, negatively impacts evaluation accuracy. Specifically, the discrepancy between the evaluator’s output and human assessments was less than 2% under our setting. On one hand, our algorithm leverages score assignments based on final steps and states, achieving a balance between computational efficiency and evaluation accuracy, on the other hand providing more context led to performance drops, significantly higher computational costs, and increased usage of evaluator LLMs, thereby straining our budget.
Figure 5: Performance correlation between automated and human evaluation across different agents, using the last X images (validated on the whole General subsets from AiTW)
In fact, empirical tests show that large models like Gemini tend to classify tasks as successful by identifying evidence of success in provided screenshots. The more information Gemini receives, the higher the probability it judges the task as successful. In few-shot scenarios with seven or eight images, adding more tokens causes token explosion, leading to hallucinations. These are reasons for our choice to limit context (instead of longer or even entire trajectories).
@@ 6.4 Training Performance
6.5 Performance Evaluation of Agents Trained with DistRL
@@ A.1 Case Study
DistRL successfully completes the task by efficiently navigating through the app’s interface. It opens the correct messaging app and enters the “New Conversation" screen without getting stuck.
@@
GPT-4V, though not specifically trained for app-specific tasks, performs well in this scenario due to its generalization capability. It opens the correct messaging app and navigates to the “New Conversation" screen successfully. GPT-4V is more flexible and suitable for simpler, general-purpose tasks. However, this general-purpose approach may not scale well for more complex tasks where app-specific expertise and interaction nuances are required.
@@
Figure 10 shows a case study on the web shopping task. DistRL demonstrates relatively smooth and fluid operations, progressing through the steps without hesitation. While it ultimately encounters early termination due to reaching the step limit (horizon), it performs each step with clear transitions and effectively navigates through the sequence. DistRL shows strong task comprehension and adaptation throughout, but its misunderstanding on the task requirement prevents it from fully completing the task. This behavior emphasizes the efficiency of DistRL’s operations and its capacity to generalize across unseen web shopping tasks, even though the task is terminated early.
@@ A.4.1 Limitations of Traditional Methods
On-policy algorithms such as Proximal Policy Optimization (PPO (Schulman et al., 2017)) and Advantage Actor-Critic (A2C (Mnih, 2016)) require synchronous data collection and updates, leading to inefficiencies in distributed and large-scale environments due to low sample efficiency and synchronization delays.
@@ A.4.2 Implementation Details
where $\alpha=0.5$ controls the extent to which prioritization is applied. A value of $\alpha=0.5$ provides a balance between uniform sampling (when $\alpha=0$) and full prioritization (when $\alpha=1$), allowing the model to benefit from both the prioritization of informative trajectories and a degree of randomness. This balance ensures that less prioritized but potentially useful experiences still have an opportunity to be replayed, helping to prevent overfitting to a narrow subset of the replay buffer.
DPER improves sample efficiency by prioritizing high-value transitions, such as those with significant temporal-difference (TD) errors or high importance sampling ratios. By focusing on these informative transitions, the agent learns more effectively from previously explored states, reducing unnecessary exploration of less relevant areas. Additionally, by maintaining policy entropy, DPER balances exploration and exploitation, accelerating convergence with fewer samples.
Policy Update Mechanism
The actor (policy network) is updated using gradients derived from the advantage estimates:
$\nabla_{\theta}\mathcal{L}=-\mathbb{E}_{\mu}\left[\rho_{t}A(s_{t},a_{t})\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})\right]-\beta\nabla_{\theta}\mathbb{E}_{\mu}\left[\log\pi_{\theta}(a_{t}|s_{t})\right]+\lambda\mathbb{E}_{\mu}\left[\mathcal{P}_{\text{invalid}}(a_{t})\right],,$ (4)
where $\theta$ represents the parameters of the policy network, $\rho_{t}$ corrects for off-policy data, $\mathbb{H}(\pi)$ encourages exploration, and $\mathcal{P}_{\text{invalid}}(a_{t})$ penalizes invalid actions. In this formula, the invalid action penalty term $\mathcal{P}_{\text{invalid}}$ penalizes actions that the agent cannot execute, addressing the challenge of entropy regularization, which encourages exploration but can lead to invalid actions. In mobile device control tasks, valid actions such as clicking and typing, while invalid commands, such as “rotate screen” in an unsupported context or interacting with non-existent UI elements, waste resources and hinder learning.
We use Gemini-1.5-Pro to evaluate each action $a_{t}$. If an action is invalid, $\mathcal{P}_{\text{invalid}}=1$; otherwise, it is 0. The penalty term is integrated into the loss as $\mathcal{L}_{\text{penalty}}=\lambda\cdot\mathbb{E}[\mathcal{P}_{\text{invalid}}]$, where $\lambda$ controls its impact. This approach ensures exploration remains within valid bounds, enhancing learning efficiency and robustness .
How is $A(s_{t},a_{t})$ calculated?
The advantage function $A(s_{t},a_{t})$ is calculated using a one-step advantage estimation approach, as follows:
$A(s_{t},a_{t})=r(s_{t},a_{t})+\gamma V(s_{t+1})-V(s_{t})$
Here, $r(s_{t},a_{t})$ comprises two components: the Monte Carlo return from timestep $t+1$ to the terminal state $s_{H}$, which is determined by success or failure signals, and hard-coded penalties for repeated actions and certain violations, serving as immediate reward signals. This combination provides a long-term cumulative reward estimate. $V(s_{t})$ represents the estimated state value at timestep $t$, $V(s_{t+1})$ is the estimated state value at the next timestep $t+1$, and $\gamma$ is the discount factor.
The trajectory value estimation function $V_{\text{traj}}$
The trajectory value function $V_{\text{traj}}$ estimates the expected return from terminal state $s_{H}$ in sparse/delayed reward environments. It serves as a baseline by labeling reward signals in collected trajectories (analogous to RLHF reward models) and filters/prioritizes trajectories for training. Through this filtering mechanism, $V_{\text{traj}}$ maintains high-quality training data, leading to more effective policy learning.
@@ A.5.1 Baseline Methods
We evaluate proprietary vision-language models (VLMs), GPT-4V(OpenAI et al., 2024) and Gemini 1.5 Pro(Reid et al., 2024), using the AppAgent framework. By applying the prompt from (Yang et al., 2023b), we enable these models to interact effectively with the environment. We assessed the AppAgent (Yang et al., 2023b) in an augmented prompting setting, where the agent explores the environment and gathers experience ahead of the inference phase. This collected experience is appended to the test-time prompt, enhancing the model’s decision-making capabilities. Unlike learning-based approaches, these methods rely on advanced prompting strategies to accomplish tasks without extensive training.
In our system’s implementation, we maintained a consistent and conflict-free approach by exclusively utilizing the T5-based Multimodal Large Language Model (MLLM) architecture (not vanilla pretrained T5, T5-Base normally has 220M, Our Agent (AutoUI-driven) has 1.3B). This encoder-decoder framework facilitates the seamless integration of various agents, enhancing the model’s versatility and performance, which is aligned with AutoUI (Zhan & Zhang, 2023) design.
Our T5-based MLLM architecture is inherently designed to support diverse decoder initializations, including the integration of pretrained decoder models. While previous works like AutoUI (Zhan & Zhang, 2023) have utilized their own model weights, we could not directly use AutoUI’s model weights to initialize our models because they were trained with AiTW knowledge. Incorporating these weights could introduce unintended biases specific to their tasks as well as prior knowledge to our training and testing data set, which might reduce the credibility and confidence of our system.
Meanwhile, our empirical analysis showed that decoder layers require more diverse initialization patterns than encoder layers to achieve optimal performance in UI-specific tasks, GPT-2 help more than the pre-trained T5 weights. Therefore, the best approach was to manually select a pretrained decoder model to initialize the weights of our T5-based decoder. We chose to initialize the decoder with GPT-2 weights, as this allowed us to leverage GPT-2’s pre-trained language generation capabilities, providing a foundational understanding of language that we could further refine through fine-tuning tailored to our specific tasks.
We did not specify the decoder type/weights in the original explanation because we found that fine-tuning efficiency was already satisfactory with GPT-2. We acknowledge that there might be better options available; however, we adopted this approach simply by drawing inspiration from our baseline DigiRL (Bai et al., 2024), which also used GPT-2 to initialize their decoders (This is a trick here). By following a structured and methodical approach from existing baselines, we ensured architectural compatibility and effective integration of the GPT-2 weights into our T5-based decoder.
@@ A.6.2 Generalization Performance on AitW Subsets
A.6.3 DistRL Agent performance on test set
Our primary contribution is improved training performance, which we evaluated using strictly separated training and testing sets. While our main results are reported in Section 6.5 using the final trained agent, we also tracked performance progression on test sets during training. Figure 13 shows the averaged results (3 trials) of periodic checkpoint evaluations on the “General" test set from AiTW, demonstrating continuous improvement throughout training.
Figure 13: Continuous Evaluation on the Test Sets during DistRL fine-tuning process
A.7 Generalizations to other domains
Expanding to a Broader Range of Mobile Applications
We conducted additional experiments on “GoogleApps" and “Install" tasks from AiTW datasets, not included in the training set, to assess generalization. Results in Appendix A.6.2 demonstrate that DistRL maintains high performance and effectively generalizes to new tasks. However, achieving high generalization remains challenging when facing substantial variations across different apps. We are also exploring more open-world benchmarks, and as future work, we plan to expand our exploration to a wider range of mobile applications.
Handling Significant UI Changes or App Updates
DistRL’s design accommodates different device types, OS versions, UI changes, and app updates by using ADB and emulators for human-like interactions and system resets. The main challenge is the compatibility of AVD and app versions. For major UI overhauls or new devices, additional training or fine-tuning may be required. However, DistRL’s architecture supports scalable data collection and training across diverse environments, enhancing its robustness in real-world scenarios. We validate this robustness using our dedicated testing infrastructure, including proprietary OS environments and hardware.
Adapting the Framework for Other OS
To extend DistRL to iOS, we can adapt our framework by replacing Android-specific tools with Apple’s counterparts (e.g., XCUITest, simctl) while maintaining our platform-agnostic core components. Our validation was facilitated by robust hardware and OS support that enabled efficient API resets and testing modules. The key challenge in migrating such RL fine-tuning platforms lies in the ability to reset states for trial-and-error learning. While our framework can support iOS environments with necessary adjustments, Android remains our primary platform due to its developer-friendly ecosystem.